# On Evaluating The Performance of Watermarked Machine-Generated Texts Under Adversarial Attacks

### v1

@@ Conclusion
Efficiency analysis revealed that both KGW and Exponential are efficient in watermark injection and detection. In terms of imperceptibility, pre-text watermarks perform better as they are embedded within token distributions and do not disrupt text fluency, unlike post-text watermarks which are more detectable due to their token manipulations. Regarding attacks, post-text attacks generally prove more efficient and practical compared to pre-text attacks, as they do not require modifications to the model’s weights. Our study also demonstrates that adversaries can significantly enhance the effectiveness of their attacks by combining different methods. This underscores the urgent need for developing more robust watermarking solutions capable of withstanding diverse and combined attack strategies.
In conclusion, while our work highlights the promise of watermarking techniques in safeguarding the authenticity of machine-generated texts, we also exposes the limitations and vulnerabilities of current approaches. Our research emphasizes the necessity for continued innovation and development of more resilient watermarking schemes to ensure the integrity and reliability of digital communications. Our code and data will be made publicly available to support further research in this domain.

### v2

@@ Conclusion
Efficiency analysis revealed that both KGW and Exponential are efficient in watermark injection and detection. In terms of imperceptibility, pre-text watermarks perform better as they are embedded within token distributions and do not disrupt text readability, unlike post-text watermarks which are more detectable due to their token manipulations. Regarding attacks, post-text attacks generally prove more efficient and practical compared to pre-text attacks, as they do not require modifications to the model’s weights. Our study also demonstrates that adversaries can significantly enhance the effectiveness of their attacks by combining different methods. This underscores the urgent need for developing more robust watermarking solutions capable of withstanding diverse and combined attack strategies.
In conclusion, while our work highlights the promise of watermarking techniques in safeguarding the authenticity of machine-generated texts, we also expose the limitations and vulnerabilities of current approaches. Our research emphasizes the necessity for continued innovation and development of more resilient watermarking schemes to ensure the integrity and reliability of digital communications. Our code and data will be made publicly available to support further research in this domain.